OpenStudy (anonymous):

Evaluate 32 + (6 - 2) • 4 - 6 over 3.

OpenStudy (anonymous):

do u know the order on how to solve expressions?

OpenStudy (uri):

PEDMAS.

OpenStudy (anonymous):

6 over 3.

OpenStudy (anonymous):

parenthesis exponents division multiplication addition subtraction

OpenStudy (anonymous):

\[\huge 32 + (6 - 2) • 4 -\frac{ 6}{ 3}\] \[\huge = 32 + 4 • 4 -2=32+ 16 -2= 48-2= 46\]